OpenStudy (anonymous):

Is sin(sin^−1x)= sin^-1(sinx) an identity? Why or why not?

OpenStudy (mimi_x3):

you mean \[\sin(\sin^{-1}x)\]?

OpenStudy (apoorvk):

NOPE - it's not sin inverse of something is always an angle. - can be from -pi/2 to pi/2 BUT sin of something is always a no. that's from -1 to 1. So, in short, never. ( i mean hardly a few times)

OpenStudy (apoorvk):

rather 'rarely'

OpenStudy (lgbasallote):

i'd hate to pop your bubble but they're both x :p try calculatoring hehe

OpenStudy (lgbasallote):

arcsine means what number should yoU "sine" to obtain a specific value...let's do numerical arcsine of 0....what will you sine to get 0? pi/2 will be the answer now we sine pi/2 which brings us back to 0 so... \[\sin (\sin^ {-1} 0) = 0\] let's try the other way around... \[\sin^{-1} (\sin 0)\] sin 0 is pi/2 now what is arcsine of pi/2? it's still 0 as welll @apoorvk :P

OpenStudy (apoorvk):

Nonsense^ arcsine of pi/2 is NOT -DEFINED - pi is 3.14 so pi/2 is something like 1.57. And in this world, sine can't be greater than one.

OpenStudy (apoorvk):

sin0 = 0. Now I know why that ":P" was for.

OpenStudy (mimi_x3):

hm.. \[\sin(\sin^{-1}x)=x\] \[\sin^{-1}(sinx) =x\] I think that this what the asker meant.

OpenStudy (apoorvk):

or girl - 'rini' sounds like a girl.

OpenStudy (lgbasallote):

*facepalm* sin 0 is 1 lol arcsine of 1 is 0 so ha :P

OpenStudy (apoorvk):

what @Mimi_x3 wrote - those two equations are separately valid - but, they are not the same things.

OpenStudy (apoorvk):

@lgbasallote you sure you are not confusing 'sine' with 'cosine'?

OpenStudy (lgbasallote):

ARGHHHH SIN 0 IS 0..>ARCSINE OF 0 IS 0..>BOTTTOMLINE THEY ARE EQUAL :PPPP

OpenStudy (anonymous):

I always thought they are equal but under certain restriction

OpenStudy (amistre64):

arcsin(sin(pi/3)) = 1.0472 sin(arcsin(pi/3)) = 1.5708 so no

OpenStudy (amistre64):

i think i forgot something on the calculator tho lol

OpenStudy (lgbasallote):

you cant arcsin pi/3...

OpenStudy (apoorvk):

pi/3 = 1.4 something.. can't be the value of a sine function

OpenStudy (amistre64):

arcsin(sin(7pi/3)) = pi/3 sin(arcsin(7pi/3)) = 7pi/3 http://www.wolframalpha.com/input/?i=sin%28arcsin%287pi%2F3%29%29

OpenStudy (lgbasallote):

well if \(-1 \le x \le 1\) then this is an identity

OpenStudy (lgbasallote):

this is like deriving an integrand equals integrating a derivative

OpenStudy (apoorvk):

HOW DOES SIN INVERSE OF 7PI/3 EXIST?? pi = 3.14 7pi/3 = around 6.5. there is an angle for which sine is '6.5'? Amazing - didn't know that!

OpenStudy (apoorvk):

Wolf just turned into a senseless werewolf. #ProgramsWillBePrograms.

OpenStudy (lgbasallote):

suspend! no one contradicts the great amistre :P lol jkjk

OpenStudy (amistre64):

wheres my button lol

OpenStudy (amistre64):

the wolf is doing complex things i believe